The shortest side (AB) of a roof truss in the form of a right triangle is x meters long. the side (BC) is 1 m longer than (AB) and the longest side (AC) is 29 m long. form an equation for x and solve it to determine the length of the other two sides.
Answer by TimothyLamb(4379)

h = sqrt( aa + bb )
a = x
b = x + 1
h = 29
---
29 = sqrt( xx + (x + 1)^2 )
29^2 = sqrt( xx + (x + 1)^2 )^2
841 = xx + (x + 1)^2
841 = xx + xx + 2x + 1
2xx + 2x + - 840 = 0
---
the above quadratic equation is in standard form, with a=2, b=2 and c=-840
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 2 -840
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 20
x = -21
---
the negative root doesn't fit the problem statement, so use the positive root:
---
x = 20
---
answer:
a = 20 m
b = 21 m
h = 29 m
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---